OpenStudy (anonymous):

Hi my son has an algebra ? that I cant help him with . Here it is : I need to find the average speed when D=10KM ,and T=23 hours I need all the steps shown . This is not a fraction problem it's algebra. Thank You !!

OpenStudy (jamesj):

Average speed, s, is defined as the distance covered, d, divided by the time, t. So here s = d/t = 10/23 km per hour.

OpenStudy (anonymous):

i need to find it in the form of a number not a fraction

OpenStudy (jamesj):

Divide 10 by 23 with a calculator.

OpenStudy (anonymous):

i nmeed the steps on how to substitute the variable

OpenStudy (jamesj):

We have substituted all of the variables already. There are no variables in the expression 10/23. So now you just need to calculate 10 divided by 23. You have a calculator, yes?

OpenStudy (anonymous):

nvm thank you